An object is being shot from a horizontal ground at an incline angle of 30 degree with respect to the ground at a speed of 49 m/s. Find the duration in seconds that the object is above the height of 13 m. Give your answer with one decimal place.


1
Expert's answer
2022-05-17T16:04:38-0400

The equation of motion of an object is given by

"y=v_0\sin\theta t-gt^2/2\\
y=49\sin30^\circ t-9.8t^2/2\\
y=24.5t-4.9t^2"

At the height 13 m we have

"13=24.5t-4.9t^2"

Roots of this equation

"t_1={0.60\:\rm s,}\;t_2=4.40\:\rm s"

Hence, duration in seconds that the object is above the height of 13 m

"\Delta t=t_2-t_1\\
=4.40-0.60=3.8\:\rm s"